Olympic Sports has two issues of debt outstanding. One is an 8% coupon bond with a face value of $24 million, a maturity of 15 years, and a yield to maturity of 9%. The coupons are paid annually. The other bond issue has a maturity of 20 years, with coupons also paid annually, and a coupon rate of 9%. The face value of the issue is $29 million, and the issue sells for 96% of par value. The firm's tax rate is 30%.

a. What is the before-tax cost of debt for Olympic? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places.)

b. What is Olympic's after-tax cost of debt? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places.)
Answer #1

1.

=(24*PV(9%,15,-8%*100,-100)/100*9%+29*96%*RATE(20,9%*100,-96,100))/(24*PV(9%,15,-8%*100,-100)/100+29*96%)=9.25%

2.

=9.25%*(1-30%)=6.48%